# Consistency conditions for O-plane unsmearing from second-order perturbation theory

### v1

@@ 1 Introduction
Given the various challenges in constructing the more cosmologically relevant de Sitter vacua [1], the more readily available compactifications to anti-de Sitter space might serve as an interesting alternative setting to study the possibility of scale-separation. Of course this assumes that the issues of scale separation and positive spacetime curvature are independent.
@@
Even in AdS, the question of scale-separation in string compactifications involves the interplay of many considerations (see [2] for a recent review). There are various arguments placing limitations on scale-separation within the paradigm of the swampland program [3, 4]. Some arguments combine the Weak Gravity Conjecture [5] with extended supersymmetry to rule out scale-separated vacua in extended supergravity theories [6]. Other restrictions come from an application of the Swampland Distance Conjecture [7] to AdS vacua. This results in the strong AdS Distance conjecture [8], which forbids scale-separation outright, with refinements proposed in [9, 10, 11]. Other arguments are based on general features of the higher dimensional supergravity equations of motion [12], similar in spirit to arguments restricting de Sitter vacua [13, 14]. While there is no all-encompassing no-go theorem, these and other considerations restrict the corners of string theory where scale-separated AdS vacua can be found.
@@
At the same time, scale-separated AdS vacua appear among some of the earliest constructions of orientifold flux compactifications. The paradigmatic example is the DGKT construction in massive type IIA strings, [15] and its various extensions [16, 17, 18]. This construction produces a family of AdS4 flux vacua, with fully stabilized moduli, that allows parametric separation between the AdS radius and the size of the internal manifold. The scale-separated limit is simultaneously a weak-coupling, large volume limit, where the whole solution can be described in terms of 10-dimensional massive type IIA supergravity with orientifold sources. Similar solutions exhibiting scale separation have also been constructed in other dimensions and duality frames [20, 19, 21, 22].
@@
These solutions either provide workarounds for some of the assumptions in the arguments against scale-separation, or for some of the stronger conjectures, appear to be explicit counterexamples. The big caveat to the DGKT-type constructions is that they are either derived within a lower-dimensional effective field theory approach, or equivalently using a smeared approximation for the O-planes [23].
@@
The smearing of sources is a common approximation in string compactifications, generally interpreted as a truncation of the full 10-dimensional equations of motion with localized sources to the zero-modes of all the fields. Since the sources actually couple to every Kaluza-Klein mode, this truncation is not a consistent one. The hope, instead, is that the truncation to zero-modes, serves as a valid approximation to the low-energy effective theory one obtains by integrating out the Kaluza-Klein modes. From the 10-dimensional perspective this equivalent to the smeared solution accurately capturing the behavior of the fields away from the sources.
@@
In practice, the relationship between smeared solutions and those with localized sources is not always straightforward. In the simplest cases of parallel BPS sources, one can construct both smeared and localized solutions and find that they approach each other in the large volume limit [24]. On the other hand, there are also examples where both smeared and localized solutions are available, but do not share the same properties [25, 26, 27]. Non-BPS sources are also known to be problematic for smearing [28].
@@
In the case of DGKT vacua, only the smeared solutions are known explicitly [23]. The construction of a full solution with localized sources is complicated by the presence of multiple source orientations, including intersection points in some limits. While not necessarily a conceptual problem, it greatly complicates the construction of explicit metrics, with simple choices of ansatz known to fail [32]. Removing the intersections by deforming the manifold, moves the technical challenge to the background manifold itself.
@@
An important step toward putting these vacua on a solid footing was the construction of leading order perturbative solutions approximating the unsmeared solution as a correction to the smeared one [29, 30]. The procedure has also been successfully applied to the lower-dimensional and massless type IIA cases [31, 20]. The success of these calculations appears to suggest a one-to-one correspondence between smeared and unsmeared DGKT-type solutions.
@@
In section 2 we motivate our investigation of the higher-order equations. We outline the sorts of troubles one encounters in perturbative approaches to non-linear problems and in gravity theories on compact spaces more specifically. We will argue for the insufficiency of the first-order result as an indicator of the existence of unsmeared solution and the necessity of investigating the higher-order equations. In section 3 we review the leading order perturbative unsmearing procedure with focus on features relevant to the second-order equations. In section 4 we will write down the second order equations and derive integral constraints on the lower order solutions focusing on the warp factor and dilaton equations. The derivation of the constraint will require careful treatment of the singular near-source behavior of the leading order corrections. We also make comment on other possible constraints as well as the effect of O-plane intersections, which appears quite benign as far as our calculations are concerned.
@@ 2 Spurious solutions in first-order perturbation theory
$x=x_{0}+\epsilon x_{1}+\frac{1}{2}\epsilon x_{2}+...\ ,$ (2.3)
@@
$x_{0}=0\implies x_{1}=0\qquad\text{or}\qquad x_{0}=y_{0}\implies x_{1}=1\ ,$ (2.4)
@@
As advertised, the problem is that the first order equation is a linearization of the original problem around the intersection of two branches of solutions and the tangent vectors to these branches span a whole vector space of solutions to the linearized equation, which are themselves not tangent to a branch of solutions. The situation is quickly remedied by going to the next order in perturbation theory. At order $\epsilon^{2}$ we have
$\frac{1}{2}x_{2}(y_{0}-x_{0})+x_{1}(1-x_{1})-\frac{1}{2}x_{0}x_{2}=0\ ,$ (2.6)
@@
$x_{1}(1-x_{1})=0\ ,$ (2.7)
once again reducing the set of valid for $x_{1}$ to only those solutions that are tangent to the branches.
@@
In the context of general relativity and specifically linearized gravity, this problem is well known, and is sometimes referred to as linearization instability [33, 34]. In particular, it occurs when studying deformations of pure GR solutions on spatially compact manifolds with killing vectors [35, 36]. A paradigmatic example is Einstein gravity on ×$\mathbb{R}\times T^{3}$. Obviously, the flat metric is a solution, differing from Minkowski space only by the periodic conditions on the spatial directions. To look for perturbative deformations, we can expand
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+\epsilon h^{(1)}_{\mu\nu}+\frac{1}{2}\epsilon^{2}h^{(2)}_{\mu\nu}+...\ ,$ (2.8)
@@
$\Box\tilde{h}_{\mu\nu}=0\ ,\qquad\tilde{h}_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}g_{\mu\nu}h^{\rho}_{~{}\rho}\ ,$ (2.9)
@@
$0=\mathcal{A}(\nabla\nabla h^{(2)})+\mathcal{B}(h^{(1)}\nabla^{2}h^{(1)})+\mathcal{C}(\nabla h^{(1)}\nabla h^{(1)})\ ,$ (2.10)
@@
More recently, examples of this linearization instability involving AdS space, rather than compact spatial manifolds, as well as some matter-coupled gravity theories have also been studied [37, 38]. In all these contexts, the conditions coming from second order perturbation theory take the form of integral constraints where a certain topological quantity must be canceled by the integral of non-linear combinations of the first-order perturbations.
@@
$\tau\sim n^{3/4}\ ,\quad w\sim n^{3/4}\ ,\quad F_{p}\sim n^{p/2}\ ,\quad H_{3}\sim n^{0}\ ,$ (2.11)
@@ 3 Scale-separated AdS and leading order unsmearing
In this section we review the relevant features of the DGKT setup and the unsmearing procedure at leading order, following the approach of [29] . The setting is massive type IIA supergravity with negative tension sources (O-planes). The metric ansatz is
$ds^{2}=g_{ij}dx^{i}dx^{j}+g_{mn}dy^{m}dy^{n}\ ,\qquad g_{ij}=w(y)^{2}\hat{g}_{ij}\ ,$ (3.1)
@@
$\rho^{(i)}=\frac{\delta^{(3)}(y_{\perp}^{(i)})}{\sqrt{g^{(i)}_{\perp}}}\ ,$ (3.4)
@@
$\rho^{(i)}_{3}=\delta^{(3)}(y^{(i)}_{\perp})dy_{\perp}^{(i)1}\wedge dy_{\perp}^{(i)2}\wedge dy_{\perp}^{(i)3}\ .$ (3.6)
@@ 3.1 Equations of motion
$d(\tau^{2}\star H_{3})+\star F_{2}F_{0}+\star F_{4}\wedge F_{2}+\star F_{6}\wedge F_{4}=0\ .$ (3.8)
@@
$dF_{4}$ $=H_{3}\wedge F_{2}\ ,\quad dF_{6}=d\star F_{6}=0\ ,\quad dH_{3}=0$ (3.9)
@@
$\mathcal{E}_{\tau}$ $=\left(\tau\frac{\delta}{\delta\tau}+\frac{1}{2}g^{MN}\frac{\delta}{\delta g^{MN}}\right)S_{IIA}$ (3.10)
@@
$\mathcal{E}_{\tau}$ $=\tau\Box\tau+\partial_{m}\tau\partial^{m}\tau+4\frac{\tau}{w}\partial_{m}\tau\partial^{m}w-\sum_{p}\frac{p-5}{4}|F_{p}|^{2}-\frac{1}{2}\tau^{2}|H_{3}|^{2}-\frac{3}{4}\tau\mu\sum_{i}\rho^{(i)}$ (3.11)
@@ 3.2 Field expansions and leading order unsmearing equations
$X=n^{\alpha_{X}}\sum_{k=0}^{\infty}\frac{1}{k!}X^{(k)}n^{-\beta_{X}k}\ ,$ (3.12)
for each field $X$, where $\alpha_{X}$ and $\beta_{X}$ are suitably chosen constants and $X^{(k)}$ are dynamical. [^2] The factor of $\frac{1}{k!}$ in the expansion is chosen to coincide with the conventions of the xPert module of the xAct package for Wolfram Mathematica [41], which was used to carry out all the perturbative expansions. The $\alpha_{X}$ are chosen to coincide with the scalings of the smeared DGKT-type solutions, i.e.
$\alpha_{g_{mn}}=\frac{1}{2}\ ,\quad\alpha_{w}=\frac{3}{4}\ ,\quad\alpha_{\tau}=\frac{3}{4}\ ,\quad\alpha_{F_{p}}=\frac{p}{2}\ ,\quad\alpha_{H_{3}}=0\ .$ (3.13)
Note that the difference in scaling between $g_{mn}$ and $w^{2}$ is ultimately responsible for the scale-separation at large $n$. The choice of $\beta_{X}$ are in principle arbitrary, but in practice are inspired by the behavior of the equations at subleading order.
@@
00 0 $=\tau^{(0)}\Box\tau^{(0)}+\partial_{m}\tau\partial^{m}\tau+4\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(0)}\partial^{m}\tau^{(0)}$ (3.15)
@@
00 0 $=\tau^{(0)2}R^{(0)}_{mn}-4\tau^{(0)2}\frac{\nabla_{m}\partial_{n}w^{(0)}}{w^{(0)}}-2\tau^{(0)}\nabla_{m}\partial_{n}\tau^{(0)}+2\partial_{m}\tau^{(0)}\partial^{m}\tau^{(0)}\ ,$
@@
We then choose the $\beta_{X}$ so that the terms that did not appear at leading order begin to contribute. This results in the values used in [29].
$\beta_{g_{mn}}=\beta_{w}=\beta_{\tau}=1$ (3.16)
@@
Other options for the scalings of subleading terms were explored in [39, 40], which allow for deformations of the original solutions at orders before the source terms enter the equations. With our current choices, at next order, we obtain for the $F_{2}$ flux
$dF^{(1)}=F_{0}H^{(0)}_{3}-\sum_{i}\mu\rho^{(i)}_{3}\ .$ (3.17)
It is at this order that the source terms first enter. In particular, integrating this equation over the internal manifold imposes tadpole cancellation, which sets $H_{3}^{(0)}$ to its value in the smeared solution. The term involving only the leading order fields can be replaced by a set of smeared source densities denoted $j^{(i)}$ and write
$dF_{2}^{(1)}=-\mu\sum_{i}(\rho^{(i)}_{3}-j^{(i)}_{3})\ ,$ (3.18)
@@
$\tau^{(0)2}d\star H_{3}^{(1)}+F_{0}\star F_{2}^{(0)}+\star F_{4}^{(0)}\wedge F_{2}^{(0)}+\star F_{6}^{(0)}\wedge F_{4}^{(0)}=0\ ,$ (3.19)
@@
$R^{(1)}_{mn}$ $\equiv\frac{1}{2}(\nabla_{p}\nabla_{m}g^{(1)~{}p}_{~{}~{}n}+\nabla_{p}\nabla_{n}g^{(1)~{}p}_{~{}~{}m}-\nabla^{m}\nabla^{n}g^{(1)~{}}_{~{}mn}-\Box g^{(1)}_{~{}mn})\ .$ (3.21)
@@
$4\frac{\Box w^{(1)}}{w^{(0)}}=-\frac{4}{3}\frac{\Box\tau^{(1)}}{\tau^{(0)}}=-2\left(\nabla^{m}\nabla^{n}g^{(1)}_{mn}-\Box g^{(1)m}_{~{}~{}~{}~{}m}\right)=\Box\Phi\ .$ (3.22)
@@
Similarly, the correction to the internal metric can be expressed as a sum over corrections sourced by each source individually. In the particular case of the $T^{6}/\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ orientifold, by choosing a traceless diagonal metric ansatz for each individual source, all the first order corrections can then be expressed in terms of the single function $\Phi$, up to integration constants [29].

### v2

# Università di Padova, Via Marzolo 8, 35131 Padova, Italy

@@ 1 Introduction
Given the various challenges in constructing the more cosmologically relevant de Sitter vacua Danielsson:2018ztv , the more readily available compactifications to anti-de Sitter space might serve as an interesting alternative setting to study the possibility of scale-separation. Of course this assumes that the issues of scale separation and positive spacetime curvature are independent.
@@
Even in AdS, the question of scale-separation in string compactifications involves the interplay of many considerations (see Coudarchet:2023mfs for a recent review). There are various arguments placing limitations on scale-separation within the paradigm of the swampland program Palti:2019pca ; vanBeest:2021lhn . Some arguments combine the Weak Gravity Conjecture Arkani-Hamed:2006emk with extended supersymmetry to rule out scale-separated vacua in extended supergravity theories Cribiori:2022trc . Other restrictions come from an application of the Swampland Distance Conjecture Ooguri:2006in to AdS vacua. This results in the strong AdS Distance conjecture Lust:2019zwm , which forbids scale-separation outright, with refinements proposed in Buratti:2020kda ; Moritz:2017xto ; Blumenhagen:2019vgj . Other arguments are based on general features of the higher dimensional supergravity equations of motion Gautason:2015tig , similar in spirit to arguments restricting de Sitter vacua Gibbons:2003gb ; Maldacena:2000mw . While there is no all-encompassing no-go theorem, these and other considerations restrict the corners of string theory where scale-separated AdS vacua can be found.
@@
At the same time, scale-separated AdS vacua appear among some of the earliest constructions of orientifold flux compactifications. The paradigmatic example is the DGKT construction in massive type IIA strings, DeWolfe:2005uu and its various extensions Camara:2005dc ; Narayan:2010em ; Marchesano:2019hfb . This construction produces a family of AdS4 flux vacua, with fully stabilized moduli, that allows parametric separation between the AdS radius and the size of the internal manifold. The scale-separated limit is simultaneously a weak-coupling, large volume limit, where the whole solution can be described in terms of 10-dimensional massive type IIA supergravity with orientifold sources. Similar solutions exhibiting scale separation have also been constructed in other dimensions and duality frames Cribiori:2021djm ; Farakos:2020phe ; Farakos:2023wps ; VanHemelryck:2022ynr ; Carrasco:2023hta .
@@
These solutions either provide workarounds for some of the assumptions in the arguments against scale-separation, or for some of the stronger conjectures, appear to be explicit counterexamples. The big caveat to the DGKT-type constructions is that they are either derived within a lower-dimensional effective field theory approach, or equivalently using a smeared approximation for the O-planes Acharya:2006ne .
@@
The smearing of sources is a common approximation in string compactifications, generally interpreted as a truncation of the full 10-dimensional equations of motion with localized sources to the internal zero-modes of all the fields. Since the sources actually couple to every Kaluza-Klein mode, this truncation is not a consistent one. The hope, instead, is that the truncation to zero-modes, serves as a valid approximation to the low-energy effective theory one obtains by integrating out the Kaluza-Klein modes. From the 10-dimensional perspective this is equivalent to the smeared solution accurately capturing the behavior of the fields away from the sources.
@@
In practice, the relationship between smeared solutions and those with localized sources is not always straightforward. In the simplest cases of parallel BPS sources, one can construct both smeared and localized solutions and find that they approach each other in the large volume limit Baines:2020dmu . On the other hand, there are also examples where both smeared and localized solutions are available, but do not share the same properties Blaback:2011nz ; Blaback:2011pn ; Apruzzi:2015zna . Non-BPS sources are also known to be problematic for smearing Blaback:2010sj .
@@
In the case of DGKT vacua, only the smeared solutions are known explicitly Acharya:2006ne . The construction of a full solution with localized sources is complicated by the presence of multiple source orientations, including intersection points in some limits. While not necessarily a conceptual problem, intersecting sources greatly complicate the construction of explicit metrics, with simple choices of ansatz known to fail in certain cases Bardzell:2024anh . Removing the intersections by deforming the manifold, when this is possible, moves the technical challenge to the background manifold itself.
@@
An important step toward putting these vacua on a solid footing was the construction of leading order perturbative solutions approximating the unsmeared solution as a correction to the smeared one Junghans:2020acz ; Marchesano:2020qvg . The procedure has also been successfully applied to the lower-dimensional and massless type IIA cases Emelin:2022cac ; Cribiori:2021djm . The success of these calculations appears to suggest a one-to-one correspondence between smeared and unsmeared DGKT-type solutions.
@@
In section 2 we motivate our investigation of the higher-order equations. We outline the sorts of troubles one encounters in perturbative approaches to non-linear problems and in gravity theories on compact spaces more specifically. We will argue for the insufficiency of the first-order result as an indicator of the existence of unsmeared solution and the necessity of investigating the higher-order equations. In section 3 we review the leading order perturbative unsmearing procedure with focus on features relevant to the second-order equations. In section 4 we will write down the second order equations and derive integral constraints on the lower order solutions focusing on the warp factor and dilaton equations. The derivation of the constraint will require careful treatment of the singular near-source behavior of the leading order corrections. We also comment on other possible constraints as well as the effect of O-plane intersections, which appears quite benign as far as our calculations are concerned.
@@ 2 Spurious solutions in first-order perturbation theory
$x=x_{0}+\epsilon x_{1}+\frac{1}{2}\epsilon^{2}x_{2}+...\ ,$ (3)
@@
$x_{0}=0\implies x_{1}=0\qquad\text{or}\qquad x_{0}=y_{0}\implies x_{1}=1\ ,$ (4)
@@
As advertised, the problem is that the first order equation is a linearization of the original problem around the intersection of two branches of solutions and the tangent vectors to these branches span a whole vector space of solutions to the linearized equation, which are themselves not tangent to any branch of solutions. The situation is quickly remedied by going to the next order in perturbation theory. At order $\epsilon^{2}$ we have
$\frac{1}{2}x_{2}(y_{0}-x_{0})+x_{1}(1-x_{1})-\frac{1}{2}x_{0}x_{2}=0\ ,$ (6)
@@
$x_{1}(1-x_{1})=0\ ,$ (7)
once again reducing the set of values for $x_{1}$ to only those solutions that are tangent to the branches.
@@
In the context of general relativity and specifically linearized gravity, this problem is well known, and is sometimes referred to as linearization instability AltasKiraci:2018rvh ; Altas:2019qcv . In particular, it occurs when studying deformations of pure GR solutions on spatially compact manifolds with killing vectors Fischer ; Deser:1973zzb . A paradigmatic example is Einstein gravity on ×$\mathbb{R}\times T^{3}$. Obviously, the flat metric is a solution, differing from Minkowski space only by the periodic conditions on the spatial directions. To look for perturbative deformations, we can expand
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+\epsilon h^{(1)}_{\mu\nu}+\frac{1}{2}\epsilon^{2}h^{(2)}_{\mu\nu}+...\ ,$ (8)
@@
$\Box\tilde{h}_{\mu\nu}=0\ ,\qquad\tilde{h}_{\mu\nu}=h_{\mu\nu}^{(1)}-\frac{1}{2}g_{\mu\nu}h^{(1)\rho}_{~{}~{}~{}\rho}\ ,$ (9)
@@
$0=\mathcal{A}(\nabla\nabla h^{(2)})+\mathcal{B}(h^{(1)}\nabla^{2}h^{(1)})+\mathcal{C}(\nabla h^{(1)}\nabla h^{(1)})\ ,$ (10)
@@
More recently, examples of this linearization instability involving AdS space, rather than compact spatial manifolds, as well as some matter-coupled gravity theories have also been studied Altas:2017fcp ; Altas:2018dci . In all these contexts, the conditions coming from second order perturbation theory take the form of integral constraints where a certain topological quantity must be canceled by the integral of non-linear combinations of the first-order perturbations.
@@
$\tau\sim n^{3/4}\ ,\quad w\sim n^{3/4}\ ,\quad F_{p}\sim n^{p/4}\ ,\quad H_{3}\sim n^{0}\ ,$ (11)
@@ 3 Scale-separated AdS and leading order unsmearing
In this section we review the relevant features of the DGKT setup and the unsmearing procedure at leading order, following the approach of Junghans:2020acz . The setting is massive type IIA supergravity with negative tension sources (O-planes). The metric ansatz is
$ds^{2}=g_{ij}dx^{i}dx^{j}+g_{mn}dy^{m}dy^{n}\ ,\qquad g_{ij}=w(y)^{2}\hat{g}_{ij}\ ,$ (12)
@@
$\rho^{(i)}=\frac{\delta^{(3)}(y_{\perp}^{(i)})}{\sqrt{g^{(i)}_{\perp}}}\ ,$ (15)
@@
$\rho^{(i)}_{3}=\delta^{(3)}(y^{(i)}_{\perp})dy_{\perp}^{(i)1}\wedge dy_{\perp}^{(i)2}\wedge dy_{\perp}^{(i)3}\ .$ (17)
@@ 3.1 Equations of motion
$d(\tau^{2}\star H_{3})+\star F_{2}F_{0}+\star F_{4}\wedge F_{2}+\star F_{6}\wedge F_{4}=0\ .$ (19)
@@
$dF_{4}$ $=H_{3}\wedge F_{2}\ ,\quad dF_{6}=d\star F_{6}=0\ ,\quad dH_{3}=0$ (20)
@@
$\mathcal{E}_{\tau}$ $=\left(\tau\frac{\delta}{\delta\tau}+\frac{1}{2}g^{MN}\frac{\delta}{\delta g^{MN}}\right)S_{IIA}$ (21)
@@
$\mathcal{E}_{\tau}$ $=\tau\Box\tau+\partial_{m}\tau\partial^{m}\tau+4\frac{\tau}{w}\partial_{m}\tau\partial^{m}w-\sum_{p}\frac{p-5}{4}|F_{p}|^{2}-\frac{1}{2}\tau^{2}|H_{3}|^{2}-\frac{3}{4}\tau\mu\sum_{i}\rho^{(i)}$ (22)
@@ 3.2 Field expansions and leading order unsmearing equations
$X=n^{\alpha_{X}}\sum_{k=0}^{\infty}\frac{1}{k!}X^{(k)}n^{-\beta_{X}k}\ ,$ (23)
for each field $X$, where $\alpha_{X}$ and $\beta_{X}$ are suitably chosen constants and $X^{(k)}$ are dynamical. [^2] The factor of $\frac{1}{k!}$ in the expansion is chosen to coincide with the conventions of the xPert module of the xAct package for Wolfram Mathematica Brizuela:2008ra , which was used to carry out all the perturbative expansions. The $\alpha_{X}$ are chosen to coincide with the scalings of the smeared DGKT-type solutions, i.e.
$\alpha_{g_{mn}}=\frac{1}{2}\ ,\quad\alpha_{w}=\frac{3}{4}\ ,\quad\alpha_{\tau}=\frac{3}{4}\ ,\quad\alpha_{F_{p}}=\frac{p}{4}\ ,\quad\alpha_{H_{3}}=0\ .$ (24)
Note that the difference in scaling between $g_{mn}$ and $w^{2}$ is ultimately responsible for the scale-separation at large $n$. The choices of $\beta_{X}$ are in principle arbitrary, but in practice are inspired by the behavior of the equations at subleading order.
@@
00 0 $=\tau^{(0)}\Box\tau^{(0)}+\partial_{m}\tau^{(0)}\partial^{m}\tau^{(0)}+4\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(0)}\partial^{m}w^{(0)}$ (26)
@@
00 0 $=\tau^{(0)2}R^{(0)}_{mn}-4\tau^{(0)2}\frac{\nabla_{m}\partial_{n}w^{(0)}}{w^{(0)}}-2\tau^{(0)}\nabla_{m}\partial_{n}\tau^{(0)}+2\partial_{m}\tau^{(0)}\partial_{n}\tau^{(0)}\ ,$
@@
We then choose the $\beta_{X}$ so that the terms that did not appear at leading order begin to contribute. This results in the values used in Junghans:2020acz .
$\beta_{g_{mn}}=\beta_{w}=\beta_{\tau}=1$ (27)
@@
Other options for the scalings of subleading terms were explored in Andriot:2023fss ; Tringas:2023vzn , which allow for deformations of the original solutions at orders before the source terms enter the equations. With our current choices, at next order, we obtain for the $F_{2}$ flux
$dF^{(1)}_{2}=F_{0}H^{(0)}_{3}-\sum_{i}\mu\rho^{(i)}_{3}\ .$ (28)
It is at this order that the source terms first enter. In particular, integrating this equation over the internal manifold imposes tadpole cancellation, which sets $H_{3}^{(0)}$ to its value in the smeared solution. The term involving only the leading order fields can be replaced by a set of smeared source densities denoted $j^{(i)}$ resulting in
$dF_{2}^{(1)}=-\mu\sum_{i}(\rho^{(i)}_{3}-j^{(i)}_{3})\ ,$ (29)
@@
$\tau^{(0)2}d\star H_{3}^{(1)}+F_{0}\star F_{2}^{(0)}+\star F_{4}^{(0)}\wedge F_{2}^{(0)}+\star F_{6}^{(0)}\wedge F_{4}^{(0)}=0\ ,$ (30)
@@
$R^{(1)}_{mn}$ $\equiv\frac{1}{2}(\nabla_{p}\nabla_{m}g^{(1)~{}p}_{~{}~{}n}+\nabla_{p}\nabla_{n}g^{(1)~{}p}_{~{}~{}m}-\nabla^{m}\nabla^{n}g^{(1)~{}}_{~{}mn}-\Box g^{(1)}_{~{}mn})\ .$ (32)
@@
$4\frac{\Box w^{(1)}}{w^{(0)}}=-\frac{4}{3}\frac{\Box\tau^{(1)}}{\tau^{(0)}}=-2\left(\nabla^{m}\nabla^{n}g^{(1)}_{mn}-\Box g^{(1)m}_{~{}~{}~{}~{}m}\right)=\Box\Phi\ .$ (33)
@@
Similarly, the correction to the internal metric can be expressed as a sum over corrections sourced by each source individually. In the particular case of the $T^{6}/\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ orientifold, by choosing a traceless diagonal metric ansatz for each individual source, all the first order corrections can then be expressed in terms of the single function $\Phi$, up to integration constants Junghans:2020acz .